Find the divergence of the following vector feild.

(i) 3xi+(y-3)j+(2-z)

(ii) Zxi+Xyj+Yzk


1
Expert's answer
2021-11-14T17:11:11-0500

The divergence of the vector feild "\bf A" is defined as

"{\rm div} {\bf A}=\frac{\partial A_x}{\partial x}+\frac{\partial A_y}{\partial y}+\frac{\partial A_z}{\partial z}"

Hence

(i)


"{\rm div} {\bf A}=\frac{\partial (3x)}{\partial x}+\frac{\partial (y-3)}{\partial y}+\frac{\partial (2-z)}{\partial z}=3"

(ii)

"{\rm div} {\bf A}=\frac{\partial (zx)}{\partial x}+\frac{\partial (xy)}{\partial y}+\frac{\partial (yz)}{\partial z}=z+x+y"